OpenStudy (amtran_bus):

If f(x)=2x-1/3 and f^-1 is the inverse of F, what is the value of f^-1(3)

OpenStudy (anonymous):

about 1.67

OpenStudy (anonymous):

because f^-1(x)=(1/2)x+1/6

OpenStudy (jdoe0001):

well, what's the \(f^{-1}(x) \ of \ 2x-\frac{1}{3} ? \)

OpenStudy (amtran_bus):

How do you get 1/3?

OpenStudy (anonymous):

2^-2(3)-1/3 2^-6-1/3 do you want it as a fraction if so this is it unless you want a decimal

OpenStudy (jdoe0001):

1/3? do you know how to find the \(f^{-1}?\)

OpenStudy (amtran_bus):

@jdoe0001 , I need refreshing on the whole thing.

OpenStudy (amtran_bus):

and I need it in whole number form. I have the answer, I just need to know how to get it. The answer is 5.

OpenStudy (jdoe0001):

if the function is f(x) = y y = ax+b then the inverse is obtained by doing a switcharoo on the variables position, so it becomes x = ay+b then you solve for "y" and is x = ay+b => x-b = ay => (x-b)/a = y so the inverse for y= ax+b would be \(\large {y = \frac{x-b}{a}}\)

OpenStudy (anonymous):

\[\frac{ 1 }{2 }x+\frac{ 1 }{ 6 } = f ^{-1}(x)\]

OpenStudy (anonymous):

\[which is = \frac{ x + \frac{ 1 }{ 3 } }{ 2 }\]

OpenStudy (amtran_bus):

I'm at this point. x=2y-1/3

OpenStudy (anonymous):

ok good, simply add the 1/3, than divide the 2

OpenStudy (anonymous):

so you will have (x+1/3)/2=f^-1(x)

OpenStudy (amtran_bus):

\[x=\frac{ 2y-1 }{ 3 }\] Where do you get subtract 1/3?

OpenStudy (amtran_bus):

I see the neg1, but you dont put the 2y over the 3 also?

OpenStudy (anonymous):

wait ooooo, thats different

OpenStudy (jdoe0001):

$$ \color{blue}{y} = 2\color{red}{x}-\frac{1}{3}\\ \text{doing a switcharoo}\\ \color{red}{x} = 2\color{blue}{y}-\frac{1}{3}\\ \text{solving for "y"}\\ \cfrac{\color{red}{x}-\frac{1}{3}}{2}=\color{blue}{y} = f^{-1}(x) $$

OpenStudy (anonymous):

no jdoe look, he put it differntly now

OpenStudy (jdoe0001):

wait... a sec

OpenStudy (jdoe0001):

$$ \color{blue}{y} = 2\color{red}{x}-\frac{1}{3}\\ \text{doing a switcharoo}\\ \color{red}{x} = 2\color{blue}{y}-\frac{1}{3}\\ \text{solving for "y"}\\ \cfrac{\color{red}{x}+\frac{1}{3}}{2}=\color{blue}{y} = f^{-1}(x) $$

OpenStudy (amtran_bus):

Here is the ORIGINAL problem: \[f(x)=\frac{ 2x-1 }{ 3}\] FInd \[f ^{-1}(3)\]

OpenStudy (anonymous):

\[\frac{ 3x+1 }{ 2 } = f^1(x)\]

OpenStudy (jdoe0001):

hmm, that looks different some :/

OpenStudy (anonymous):

f^-1(3)=5

OpenStudy (anonymous):

no wonder we did it wrong, you gave it to us wrong.

OpenStudy (amtran_bus):

That is real nice. :(

OpenStudy (anonymous):

so yes the answer is 5.

OpenStudy (amtran_bus):

Can you work it out step by step? (patience please)

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

$$ y=\frac{ 2x-1 }{ 3}\\ \text{doing a switcharoo}\\ x=\frac{ 2y-1 }{ 3}\\\\ \text{solving for "y"}\\ $$

OpenStudy (jdoe0001):

whatever "y" gives you, is your \(f^{-1}(x)\)

OpenStudy (amtran_bus):

Thanks, got it!